### v2

# Minimal length: A source of quantum non-locality

Abstract
The narrow and subtle difference between the Hilbert spaces of operators corresponding to the purely quantum mechanical momentum and the generalized momentum that includes minimal length effects is polished. Additionally, the existence of complex numbers in quantum mechanics may be justifiable as a consequence of minimal length. A novel quantum entanglement generation is also reported indicating the power of theories including a minimal length in enriching the current understanding of quantum non-locality.
I Introduction
Based on the Heisenberg Uncertainty Principle (HUP), there is no limit on the minimum value of length, provided that the momentum information is completely lost. On the other hand, the role of gravity at quantum scale is mysterious Hossenfelder:2012jw ; Das:2008kaa ; Ali:2009zq ; Ali:2011fa . Nevertheless, considering Newtonian gravity in revisiting the Heisenberg microscope unveils a non-zero minimal length Hossenfelder:2012jw ; Mead:1964zz ; Mead:1966zz . It leads to the Generalized Uncertainty Principle (GUP), which is also reinforced in quantum scenarios of gravity Hossenfelder:2012jw ; Kempf:1994su . HUP lies at the heart of classically amazing phenomena like quantum non-locality Einstein:1935rr and correspondingly, the existence of minimal length and quantum aspects of gravity affect the quantum entanglement, the violation of Bell’s inequality Bose:2017nin ; Marletto:2017kzi ; Blado:2017gnf ; Aghababaei:2021yzx ; Aghababaei:2022rqi ; Park:2022jrs ; Moradpour:2023gsn ; Filho:2023vpx ; Marletto:2025fpm , and the information bounds Rastegin:2016xeb ; Pedram:2016gps ; Aghababaei:2022jxd . Indeed, we are even witnessing a fascinating claim that classical gravity produces quantum entanglement Aziz:2025ypo ; Marletto:2025fpm . Additionally, the conjectured correspondence between metric-affine gravity and quantum gravity scenarios Ali:2024tbd further motivates us to study the implications of the minimum length on various phenomena such as quantum non-locality and related topics.
In Quantum Mechanics (QM), momentum ($\hat{p}$) and position operator ($\hat{x}$) are conjugate and thus Fourier transform dual. They do not commute with each other (the origin of the EPR paradox), and it is enough to replace momentum and position in a classical Hamiltonian with the corresponding operators to get the quantum counterpart of the system. Usually, the position representation is taken into account to write the Schrödinger equation for which $\hat{p}=-i\hbar\nabla$ Brau:1999uv . In this manner, the effects of minimal length are gathered in the change of momentum operator from canonical momentum ($\hat{p}$) to generalized momentum ($\hat{P}$) as $\hat{p}\rightarrow\hat{P}=\hat{p}~(1+\beta\hat{p}^{2})$, up to the first order of the GUP parameter $\beta$ Hossenfelder:2012jw ; Kempf:1994su ; Brau:1999uv . Correspondingly, to study the implications of minimal length on a quantum mechanical system with Hamiltonian $\hat{H}$, it is enough to replace $\hat{p}$ with $\hat{P}$ Hossenfelder:2012jw ; Kempf:1994su . Briefly, while $\hat{P}$ is claimed to be the physical momentum operator in the presence of minimal length, canonical momentum $\hat{p}$ is still an operator with its eigenvalues and Hilbert space.
The EPR intelligently experiment Einstein:1935rr reveals the non-local nature of QM that finally opens a new window towards information science by relying on quantum non-locality capacity, originally rooted in HUP Einstein:1935rr ; Bohm:1957zz ; Heisenberg ; Franson ; oppen ; Peres:2002ip . Despite previous works on $i$) testing the quantum nature of gravity using entangled states Bose:2017nin ; Marletto:2017kzi ; Altamirano:2016fas ; Hall:2017nzl ; Doner:2022npz ; Spaventa:2023eys and $ii$) the effects of GUP on Bell inequality, quantum non-locality, and thus information bounds Blado:2017gnf ; Aghababaei:2021yzx ; Aghababaei:2022rqi ; Park:2022jrs ; Moradpour:2023gsn ; Filho:2023vpx ; Pedram:2016gps ; Rastegin:2016xeb ; Aghababaei:2022jxd , the role and ability of minimal length (GUP or quantum aspects of gravity) to directly generate quantum non-locality and entangled states have not yet been studied.
In summary, the commutator $\hat{x},\hat{p}$] is the cornerstone of the EPR argument, and the existence of minimal length (GUP) replaces it with $\hat{x},\hat{P}$]. Additionally, the existence of complex numbers in QM is a challenging issue AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi that is also predicted to influence Bell-type experiments in favor of the complex Hilbert-space formulation of QM Renou:2021dvp . Hence, on one hand, GUP (the existence of minimal length) affects quantum entangled states and Bell’s inequality violation criteria, Aghababaei:2021yzx ; Aghababaei:2022rqi ; Moradpour:2023gsn . On the other hand, Bell-type experiments (quantum non-locality) seem able to verify the existence of complex numbers in QM Renou:2021dvp . Consequently, two questions need to be answered: Does minimal length $i$) support complex numbers in QM? and $ii$) Is it a source of quantum non-locality?
Therefore, if entangled states are found out, provided that they $i$) originate from minimal length and $ii$) include complex quantum numbers, then both questions are answered. In the next section, the relationship between the states (spaces) of $\hat{p}$ and $\hat{P}$ is investigated, leading to a justification for considering complex eigenvalues in QM. After that, and equipped with the results of the previous section, the ability of minimal length, leading to GUP, in generating quantum non-locality is studied by providing entangled states that also include complex quantum numbers. A summary is presented in the last section.
II Canonical momentum, generalized momentum, and complex numbers
In the purely quantum mechanical regime, we basically have $\hat{x},\hat{p}]=i\hbar$ and
$\begin{split}&\hat{x}~|x_{0}\rangle=x_{0}~|x_{0}\rangle~\longrightarrow~\langle x|x^{\prime}\rangle=\delta(x-x^{\prime}),\\ &\hat{p}~|p_{0}\rangle=p_{0}~|p_{0}\rangle~\longrightarrow~\langle p\ |p^{\prime}\rangle=\delta(p-p^{\prime}),\\ &\int|p\rangle\langle p|~dp~=~\int|x\rangle\langle x|~dx=\hat{1},\\ &\langle x|\hat{p}|p\rangle=p~\langle x|p\rangle=-i\hbar\partial_{x}\langle x|p\rangle\Rightarrow\langle x|p\rangle=\frac{\exp(\frac{ipx}{\hbar})}{\sqrt{2\pi\hbar}}.\end{split}$ (1)
Therefore, $\exp(-\frac{ipx}{\hbar})$ corresponds to state $|-p\rangle$ with eigenvalue p-p. The coefficient $\frac{1}{\sqrt{2\pi\hbar}}$ has been adopted to satisfy $\langle p\ |p^{\prime}\rangle=\delta(p-p^{\prime})$.
In the presence of minimal length, the generalized momentum operator $\hat{P}(\equiv\hat{p}~(1+\beta\hat{p}^{2}))$ should be employed to measure momentum. It is also worthy to remember that, based on quantum mechanics, the canonical momentum operator ($\hat{p}$) is still a physical observable with its Hilbert space ($|p\rangle$). In the minimal length regime, the position operator remains unchanged (position representation Brau:1999uv ). We have $\hat{x},\hat{P}]=i\hbar(1+3\beta\hat{P}^{2})$ leading to Hossenfelder:2012jw
$\Delta x\Delta P=\frac{\hbar}{2}(1+3\beta(\Delta P)^{2}),$ (2)
that admits $\Delta x_{min}=\sqrt{3\beta}~\hbar\equiv\sqrt{\beta_{0}}~l_{p}$ where $l_{p}$ denotes the Planck length, respectively Hossenfelder:2012jw . In this regime, we have $\hat{H}=\frac{\hat{P}^{2}}{2m}$ for the Hamiltonian. Now, let us represent the eigenstate of $\hat{P}$ with $|P\rangle$ and thus $\hat{P}~|P_{0}\rangle=P_{0}~|P_{0}\rangle$. Moreover, since $\hat{P}$ and $\hat{p}$ commute, their info can be represented simultaneously. But, it should be noted that since $\hat{P}\neq\hat{p}$, their states (Hilbert spaces) are not the same ($|p\rangle\neq|P\rangle$), a point easily verifiable using $\langle x|P\rangle\neq\langle x|p\rangle$. Indeed,
$\langle x|\hat{P}|P\rangle=P~\langle x|P\rangle=-i\hbar\partial_{x}(1-\beta\hbar^{2}\partial_{x}^{2})\langle x|P\rangle,$ (3)
whose solutions take the form
$\exp(\frac{i\mathcal{P}x}{\hbar}),$ (4)
with these values of $\mathcal{P}$
$\mathcal{P}_{1}=\frac{\mathcal{A}}{6\beta}-\frac{2}{\mathcal{A}},$ (5)
$\mathcal{P}_{2}=-\frac{\mathcal{P}_{1}}{2}+i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg],$
$\mathcal{P}_{3}=-\frac{\mathcal{P}_{1}}{2}-i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg],$
$\mathcal{A}=\bigg(\big[108P+12\sqrt{3}\sqrt{\frac{27P^{2}\beta+4}{\beta}}\big]\beta^{2}\bigg)^{\frac{1}{3}}.$
Interestingly enough, it is seen that Eq. (4) is also the eigenstates of the canonical momentum operator ($\hat{p}$) with eigenvalues $\mathcal{P}_{i}$. Since $\beta$ is positive Hossenfelder:2012jw ; Kempf:1994su , when $\beta\rightarrow 0^{+}$, we have $\mathcal{P}_{1}\rightarrow P(\rightarrow p)$, $\mathcal{P}_{2}\rightarrow i\infty$, and $\mathcal{P}_{3}\rightarrow-i\infty$. Therefore, in the presence of minimal length, $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ have also real part ($\frac{\mathcal{P}_{1}}{2}$). In fact, whenever $\beta\neq 0$, while for each eigenvalue $p$ there is only one eigenvalue $P=p(1+\beta p^{2})$, the inverse is not necessarily always true as $p(1+\beta p^{2})=P$ is a third order equation of $p$. Hence, there is a threefold degeneracy for each value of $P$ that can be unveiled using $\hat{p}$. It is worth noting that the existence of complex values of momentum and indeed, the implications of complex numbers in QM are longstanding topics AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Consequently, the existence of complex momentum in quantum mechanics may be a consequence of minimal length.
$|P\rangle$ and $|p\rangle$ denote the basis of the Hilbert spaces corresponding to the operators $\hat{P}$ and $\hat{p}$, respectively, and by considering the above argument, one can write the general solution of Eq. (3) as
$|P\rangle=\sum_{k=1}^{3}\alpha_{k}|p=\mathcal{P}_{k}\rangle,$ (6)
where $\alpha_{i}$ denotes the expansion coefficients, which brings us to
$\langle x|P\rangle$ $=$ $\sum_{k=1}^{3}\alpha_{k}\langle x|p=\mathcal{P}_{k}\rangle=\frac{1}{\sqrt{2\pi\hbar}}\sum_{k=1}^{3}\alpha_{k}\exp(\frac{i\mathcal{P}_{k}x}{\hbar}),$
$\langle p|P\rangle$ $=$ $\int_{-\infty}^{\infty}\langle p|x\rangle\langle x|P\rangle dx$ (7)
$=$ $\sum_{k=1}^{3}\alpha_{k}\big(\frac{1}{2\pi\hbar}\big)\int_{-\infty}^{\infty}\exp(-\frac{ipx}{\hbar})\exp(\frac{i\mathcal{P}_{k}x}{\hbar})dx$
$=$ $\sum_{k=1}^{3}\alpha_{k}\delta(p-\mathcal{P}_{k}).$
Before proceeding further, it should be noted that Kempf:1994su
$\langle P_{1}|P_{2}\rangle=\big(1+\beta P_{1}^{2}\big)\delta(P_{1}-P_{2}),$ (8)
$\hat{1}=\int|P\rangle\langle P|\frac{dP}{1+\beta P^{2}},$
and additionally, Eq. (1) is still valid for operators $\hat{x}$ and $\hat{p}$, of course, in their spaces. As the next step, since $dp=\frac{dP}{1+\beta P^{2}}$ Hossenfelder:2012jw , one can write
$\langle P_{1}|P_{2}\rangle=\sum_{k,m}\alpha_{k}^{\ast}\alpha_{m}\langle p_{k}|p_{m}\rangle\Rightarrow$ (9)
$\int\langle P_{1}|P_{2}\rangle\frac{dP_{2}}{1+\beta P_{2}^{2}}=\sum_{k,m}\int\alpha_{k}^{\ast}\alpha_{m}\langle p_{k}|p_{m}\rangle dp_{m},$
that leads to
$\sum_{k=1}^{3}|\alpha_{k}|^{2}=1.$ (10)
Clearly, if measurement is done in the $|P\rangle$ bases, then we deal with only one state and thus one outcome. On the other hand, if a purely quantum-mechanical apparatus is employed (or equivalently, the measurement is performed in the $|p\rangle$ space), then three outcomes $\mathcal{P}_{i}$ are likely to occur after the collapse of the wave function.
III Quantum entanglement production
By coming back to Eqs. (3-5), it is easily seen that $\mathcal{P}_{k}$ is a solution corresponding to state $|-P\rangle$ with eigenvalue P-P (just same as $\exp(-\frac{ipx}{\hbar})$ and p-p in the purely quantum mechanical regime (1)) and interestingly enough, we have
$-\mathcal{P}_{k}=\mathcal{P}_{l}+\mathcal{P}_{m},$ (11)
where $k,l,m\in\{1,2,3\}$. Comparing these results with Eq. (5) for which
$P=\mathcal{P}_{j}(1+\beta\mathcal{P}_{j}^{2}),$ (12)
where $j=1,2,3$, one easily reaches
$-\mathcal{P}_{j}(1+\beta\mathcal{P}_{j}^{2})=-P,$ (13)
yielding
$|-P\rangle=\sum_{k=1}^{3}\gamma_{k}|-\mathcal{P}_{k}\rangle.$ (14)
Therefore, $|-P\rangle$ consists of eigenstates of canonical momentum with eigenvalues $\mathcal{P}_{j}$.
At this step, consider a two-particle system whose total momentum in the presence of minimal length (or equally, the total generalized momentum) is zero so that $P_{1}=-P_{2}\equiv P$ and they are moving along an arbitrary direction (such as the $x$ axis). Clearly, the momentum state $|P,-P\rangle$ indicates that the outcome of measuring $P_{i}$ is completely predictable. Interestingly enough, the generalized momentum conservation ($\sum_{j=1}^{j=2}P_{j}=0$) in the $\hat{P}-$space also inspires the canonical momentum conservation ($\sum_{j=1}^{j=2}p_{j}=$0) generating the $p_{1}=-p_{2}$ condition. Accordingly, if purely quantum mechanical momentum measurements are employed (or equally, the $|p\rangle$ space is focused on), then Eqs. (6) and (14) lead us to the state
$\sum_{i=1}^{3}\alpha_{i}\gamma_{i}|\mathcal{P}_{i},-\mathcal{P}_{i}\rangle,$ (15)
apart from a normalization constant. It unveils a novel quantum non-locality that exposes itself in the canonical momentum space. Therefore, the existence of minimal length seems capable to generate quantum non-locality.
IV Summary and Concluding Remarks
The discovered degeneracy in $\langle x|P\rangle$ seems breakable using the canonical momentum operator. It also signals to us that minimal length (and thus GUP) may be an origin for the emergence of complex numbers in QM, which play impressive roles in various phenomena such as quantum non-locality AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Finally, and most importantly, calculations reveal the ability of minimal length to create quantum entanglement. The simplest spin versions of the EPR thought experiment Bohm:1957zz include a two-$1/2$ spin particle system whose spin state in the total spin space is either $|S=1,S_{z}=0\rangle$ or $|S=0,S_{z}=0\rangle$. The states’ decomposition using the spin state of each particle leads to the $\frac{1}{\sqrt{2}}\big[|+,-\rangle+|-,+\rangle\big]$ and $\frac{1}{\sqrt{2}}\big[|+,-\rangle-|-,+\rangle\big]$ states, respectively, addressing the quantum entanglement. The same behaviour is also obtained by expressing (measuring) $|P,-P\rangle$ in the canonical momentum bases (using canonical momentum operator) of each particle, leading to Eq. (15).
Acknowledgment
This work is based upon research funded by Iran National Science Foundation (INSF) under project No. 4046949.
Data availability
No data was used for the research described in the article.
Declaration of competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.